|2x + 6| - 4 = 20
Thanks i need help solving

Answer by edjones(7311)

|2x + 6| - 4 = 20
|2x+6|=24
.
2x+6=24
2x=18
x=9
.
2x+6=-24
2x=-30
x=-15
.
Ed